OpenStudy (anonymous):

A Carpenter charges a set fee for each job in a addition to his hourly charge. If a carpenter bills a customer $125 for a job which takes him 3 hours and bills a second customer $200 for a 6 hour job, how much would he charge for a job that takes him 8 hours? plzzzzz help!!!!

OpenStudy (anonymous):

set up your equations. we have two unknowns. x will be the set fee and y*t will be the hourly rate times the number of hours. can you set up from here?

OpenStudy (anonymous):

not really can u help me?

OpenStudy (anonymous):

for the first job takes 3 hours, so t=3. the total cost is 125 so, x+3y=125 try to set up the second equation using the second job's time and cost.

OpenStudy (anonymous):

ok thnx and the other 1 would be x+6y+200 am i right?

OpenStudy (anonymous):

=200

OpenStudy (anonymous):

o my bad i meant that lol

OpenStudy (anonymous):

then solve for x in one equation and so on

OpenStudy (anonymous):

hmmm how can u do it?

OpenStudy (anonymous):

take the first equation I gave you and move everything to the other side except for x. that is what x equals. plug that into the second equation for x. this leaves you with only y's and constants.

OpenStudy (anonymous):

ok let me do it hold on

OpenStudy (anonymous):

so it would be x=3y+125?

OpenStudy (anonymous):

125-3y

OpenStudy (anonymous):

121? then

OpenStudy (anonymous):

now plug (125-3y) into the second equation in place of x.

OpenStudy (anonymous):

can u help me cuz im totally lost lol

OpenStudy (anonymous):

I'm trying to show you the steps needed. so the second equation is: x+6y=200 and we know that: x=125-3y so (125-3y)+6y=200 is the same thing as the original second equation

OpenStudy (anonymous):

ok thnx and i multiply 6x with -3x right?

OpenStudy (anonymous):

now combine the y's and move constants to the right side

OpenStudy (anonymous):

-18=325? am i right

OpenStudy (anonymous):

125-3y+6y=200, I put the parenthesis in last time to show you it was x, but the -3y is not multiplied by 6y it is added to it. remember, if you move a positive to the other side, what you are doing is subtracting it from both sides.

OpenStudy (anonymous):

ok is it 3y=75?

OpenStudy (anonymous):

yes, now divide by three to solve for y. then go to the equation x=125-y to solve for x. remember that x is the set fee and y is the hourly rate.

OpenStudy (anonymous):

so the answer will be 100?

OpenStudy (anonymous):

cuz y=25

OpenStudy (anonymous):

x would equal 100

OpenStudy (anonymous):

x=100, yes. now what do you think your equation to answer the question is?

OpenStudy (anonymous):

x=100 and y=25?

OpenStudy (anonymous):

plz help me i almost gtg

OpenStudy (anonymous):

yes on values. remember what we defined x and y as are the hourly rate and the set fee. so we need to add up the set fee plus the product of the hourly rate times the hours. this will give you the total cost for the job.

OpenStudy (anonymous):

so wats the answer?

OpenStudy (anonymous):

x+8y=

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

you have x and y

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

plz help me i almost gtg

OpenStudy (anonymous):

100+8*25=the cost of the job

OpenStudy (anonymous):

300$?

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

it is.

OpenStudy (anonymous):

ok thank you!!!!!!

OpenStudy (anonymous):

np